### v1

# Strong discrete Morse theory

Abstract.
The purpose of this work is to develop a version of Forman’s discrete Morse theory for simplicial complexes, based on internal strong collapses. Classical discrete Morse theory can be viewed as a generalization of Whitehead’s collapses, where each Morse function on a simplicial complex $K$ defines a sequence of elementary internal collapses. This reduction guarantees the existence of a CW-complex that is homotopy equivalent to $K$, with cells corresponding to the critical simplices of the Morse function. However, this approach lacks an explicit combinatorial description of the attaching maps, which limits the reconstruction of the homotopy type of $K$. By restricting discrete Morse functions to those induced by total orders on the vertices, we develop a strong discrete Morse theory, generalizing the strong collapses introduced by Barmak and Minian. We show that, in this setting, the resulting reduced CW-complex is regular, enabling us to recover its homotopy type combinatorially. We also provide an algorithm to compute this reduction and apply it to obtain efficient structures for complexes in the library of triangulations by Benedetti and Lutz.
Key words and phrases:
Discrete Morse theory, strong homotopy theory, simplicial complexes, posets, collapses
2020 Mathematics Subject Classification:
55U10 (primary), 57Q05, 05E45, 68W05 (secondary)
1. Introduction
The combinatorial analysis of homotopy types dates back to the 1940s with Whitehead’s simple homotopy theory [32, 33, 34, 35]. In an effort to describe homotopy equivalences through rigid transformations of combinatorial structures, Whitehead introduced collapses and expansions for finite simplicial complexes. These moves were inspired by Tietze transformations for group presentations. While Tietze’s theory provides a complete set of transformations between equivalent presentations, Whitehead’s simple homotopy theory is more restrictive and fails to capture all homotopy deformations. Nevertheless, it remains a powerful combinatorial framework, influencing classical problems [1, 24]; such as Zeeman’s conjecture [36]; the Andrews–Curtis conjecture [2, 19]; and the Poincaré conjecture [28, 37]; as well as modern computational approaches to analysing complexes and data [4, 5, 29].
Whitehead’s elementary deformations are called collapses (and their reverse moves, expansions), which rely on the concept of a free face. If $\tau$ is the unique simplex of a simplicial complex $K$ properly containing $\sigma$, then both $\sigma$ and $\tau$ can be removed from $K$ while preserving its homotopy type and simplicial structure (see Figure 1, left). However, many simplicial complexes lack free faces (e.g., triangulations of closed manifolds). A more general concept is that of an internal free face, leading to an internal collapse [20, 22, 15]. Here, $\sigma$ becomes a free face of $\tau$ in a subcomplex $L\subseteq K$. To preserve the homotopy type of $K$ after removing $\sigma$ and $\tau$, the simplices in $K\smallsetminus L$ must be re-attached, often losing the simplicial structure (see Figure 1, right). In general, there is no explicit or computable method to determine such attachments in arbitrary dimensions.
Figure 1. Simple collapse (left) vs internal simple collapse (right).
Discrete Morse theory [17, 18, 21], developed by Forman in the 1990s as a combinatorial analogue of smooth Morse theory for manifolds, is a tool for studying the homotopy type of simplicial complexes via real-valued functions defined on them. Concretely, given a simplicial complex $K$ and a discrete Morse function $f\colon K\to\mathbb{R}$, there exists a reduced CW-complex built from the critical simplices of $f$, denoted by ${\mathrm{core}}_{f}(K)$, which is homotopy equivalent to $K$. Moreover, a recent result establishes a connection between discrete Morse theory and simple homotopy theory: if $K$ has dimension $N$, then there exists a sequence of collapses and expansions from $K$ to ${\mathrm{core}}_{f}(K)$ through intermediate complexes of dimension at most $N+1$ (also called an $(N+1)$-deformation and denoted by $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0pt{\mathrm{core}}_{f}(K)$[15]. This result relies on the fact that discrete Morse functions are equivalent to well-ordered sequences of internal collapses, which encode the deformation from $K$ to ${\mathrm{core}}_{f}(K)$ [15, 22]. The construction of the reduced complex ${\mathrm{core}}_{f}(K)$, however, often sacrifices the original simplicial and combinatorial structure of $K$. The resulting attaching maps tend to be more intricate and are not explicitly determined, transferring the topological complexity of the space from many simplices with simple attaching maps to fewer cells (corresponding to the critical simplices) whose attaching maps are more complex. This work focuses on a particular case of discrete Morse theory in which it is possible to provide an explicit and computable method for describing these attaching maps, even in higher dimensions.
In 2012, a strong version of homotopy theory was developed by Barmak and Minian [3]. This approach builds upon simple homotopy theory by introducing strong collapses of simplicial complexes. Inspired by vertex reductions on posets [31], strong collapses involve eliminating vertices $v$ (called dominated vertices) whose link is a simplicial cone. In general, removing (the open star of) vertices with a contractible link preserves the homotopy type, as a consequence of the Gluing Theorem [11]. Coned links, however, offer a systematic way to simple collapse $K$ to the subcomplex $K\smallsetminus v$ (see Figure 2, left).
Strong collapses, as collections of simultaneous collapses, are highly efficient. Unlike Whitehead’s collapses, strong collapses always yield the same minimal irreducible subcomplex — the strong core of $K$. However, the strong core of $K$ often coincides with $K$ itself, as many simplicial complexes lack dominated vertices. To overcome this limitation, one can consider performing strong collapses internally, where domination is restricted to a subcomplex of $K$ (see Figure 2, right). In such cases, a strategy must again be developed to recompute the attaching maps of the remaining cells, potentially compromising the simplicial structure of the reduced complex.
Figure 2. Strong collapse (left) vs internal strong collapse (right).
In this work, we establish a method for studying internal strong collapses and prove that the reduced CW-complex obtained from any sequence of such collapses is always regular (i.e., its attaching maps are homeomorphisms onto their images). As a consequence, the CW-structure of the reduced complex is fully combinatorially determined by the incidences of its cells. More broadly, we develop the theoretical foundations of strong Morse theory, a version of discrete Morse theory in which discrete Morse functions correspond to sequences of internal strong collapses. We propose an efficient algorithm for computing the reduced complex ${\mathrm{core}}_{f}(K)$ for such discrete Morse functions $f$. To demonstrate its applicability, we use this method to identify more efficient regular models of the simplicial complexes in the Library of Triangulations [5, 6, 7] by Benedetti and Lutz.
1.1. Motivation and related work
Classical discrete Morse theory does not provide an explicit description of the reduced homotopy equivalent CW-complex. A discrete Morse function on a simplicial complex $K$ induces a collection of critical simplices that correspond to the cells in the reduced CW-complex, along with information about their incidences, which is sufficient to determine the homology of the simplicial complex $K$. However, recovering the homotopy type of $K$ from the reduced CW-complex requires a more refined understanding of the attaching maps. Hence, except in special cases—such as when the homotopy type is fully determined by the list of critical cells (e.g., wedges of spheres)—the algorithmic reconstruction of homotopy types from discrete Morse data remains an open problem.
Some progress has been made toward addressing this challenge. In [15], the author develops a combinatorial framework to describe the attaching maps of the reduced CW-complex for 2-dimensional simplicial complexes equipped with a discrete Morse function with a single critical 0-simplex. The approach makes strong use of the correspondence between 2-complexes and group presentations. However, it does not extend to higher dimensions, as the homotopy types of $N$-dimensional complexes with $N>2$ generally cannot be fully captured combinatorially via group presentations. Rather than attempting to recover arbitrary attaching maps, the present work focuses on cases in which discrete Morse functions yield reduced complexes with particularly simple attaching maps—namely, homeomorphisms.
Our central goal is to develop a combinatorial and computable framework for a strong discrete Morse theory that allows for the explicit reconstruction of the reduced CW-complex. By restricting Morse functions to those induced by internal strong collapses, we prove that the resulting reduced complex is always a regular CW-complex, and hence its homotopy type can be entirely determined by the poset of incidence cells. Moreover, as strong Morse theory is a restricted version of classical discrete Morse theory, the reduced complex $(N+1)$-deforms to the original simplicial complex $K$, where $N$ is the dimension of $K$. Consequently, this theory also provides a computational method to perform $(N+1)$-dimensional deformations.
In [16], the authors also establish connections between discrete Morse theory and strong collapses, focusing on general discrete Morse functions. Their work introduces a classification of internal collapses into critical and regular pairs, providing criteria to determine whether a sub-sequence of internal collapses is induced by a strong collapse. By contrast, our work only focuses on the specific class of Morse functions derived from only internal strong collapses.
A complementary perspective is presented in [26, 25], where the authors encode the homotopy type of a simplicial complex endowed with a discrete Morse function into a combinatorially defined flow category. The classifying space of the flow category recovers the homotopy type of the original complex. However, these categories are often large and computationally challenging, limiting their algorithmic usage in homotopy reconstruction.
1.2. Main results and outline
Let $K$ be a finite simplicial complex, and let $g\colon V(K)\to\mathbb{R}$ be a real-valued function on its set of vertices. The function $g$ induces a filtration of $K$ by subcomplexes, where each subcomplex consists of all simplices whose vertices lie below a given threshold in $g$. For each vertex $v\in V(K)$, we consider the first subcomplex in the filtration that contains $v$, and define its descending open star as the collection of simplices in that subcomplex that contain $v$. A vertex is called strong critical if, at the filtration step where it first appears, it is not dominated by any vertex from an earlier subcomplex.
This filtration of $K$ leads to a strong homotopy analogue of the classical discrete Morse lemmas (Lemma 2.4). Intervals in the filtration without strong critical vertices correspond to strong collapses, while those with strong critical vertices introduce changes in the homotopy type, determined by the simplices in its descending open stars (Lemma 3.6). This reduction process produces a sequence of internal strong collapses, resulting in a reduced complex, called the strong internal core and denoted by ${\mathrm{core}}_{g}(K)$. Notably, we prove that this reduced CW-complex is regular, that is, its attaching maps are homeomorphisms with its image.
Our main result in strong Morse theory is as follows.
Theorem (Strong Morse Theory).
Let $K$ be a finite simplicial complex, and let $g\colon V(K)\to\mathbb{R}$ be a real-valued function on the vertex set of $K$. Then $K$ is homotopy equivalent to a regular CW-complex ${\mathrm{core}}_{g}(K)$, whose cells are in one-to-one correspondence with the simplices in the descending open stars of the strong critical vertices of $g$. Moreover, if $\dim(K)=N$, then $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0pt{\mathrm{core}}_{g}(K)$.
We also show that this result is a particular case of classical discrete Morse theory (Theorem 3.9). More precisely, for every function $g\colon V(K)\to\mathbb{R}$, there exists a discrete Morse function $f\colon K\to\mathbb{R}$ such that the critical simplices of $f$ are exactly the simplices in the descending open stars of the strong critical vertices of $g$. Moreover, the CW-complexes ${\mathrm{core}}_{f}(K)$ and ${\mathrm{core}}_{g}(K)$ coincide. [^1] By a slight abuse of notation, we denote both by ${\mathrm{core}}$ depending on the context: ${\mathrm{core}}_{f}(K)$ denotes the internal core obtained from a discrete Morse function $f$ on simplices, and ${\mathrm{core}}_{g}(K)$ denotes the strong internal core obtained from a function $g$ on vertices.
The combinatorial structure of the strong internal core admits a computational interpretation of internal strong collapses, which in turn leads to an efficient algorithm for their computation (Algorithm 2). An implementation is publicly available at https://github.com/ximenafernandez/Strong-Morse-Theory. To illustrate the effectiveness of this method, we apply it to examples from the Library of Triangulations [6], a repository of challenging simplicial complexes frequently used in the study of homotopy types. The results of the reductions are presented in Tables 1 and 2.
Beyond its role in understanding homotopy types, the algorithmic reduction of simplicial complexes also has applications in topological data analysis, where the complexity of persistent homology computations increases rapidly with input size. For instance, computing degree-$d$ persistent homology of the Vietoris–Rips filtration of a point cloud of size $N$ has worst-case complexity $O(N^{3(d+2)})$ [27]. The method introduced in [10] uses strong collapses to simplify Vietoris–Rips filtrations prior to computing persistent homology. Since our results extend algorithmic simplification of complexes via internal strong collapses, they have potential applications in further optimising such filtration-based computations.
The remainder of this manuscript is organized as follows. Section 2 describes discrete Morse theory as generalized Whitehead’s collapses. In Section 3, we study a strong version of discrete Morse theory as a generalization of strong collapses, which allows us to explicitly recover the combinatorial structure of the reduced CW-complex. Section 4 describes an algorithm for the construction of the strong internal core in terms of posets and acyclic matchings. Finally, Section 5 discusses the algorithm for random strong Morse reduction and presents experiments on the Library of Triangulations [6].
Notation. We use $\simeq$ for homotopy equivalences, $\cong$ for homeomorphisms, and ||$\cdot$|| ⋅ | for geometric realizations. The symbols $\sim$ and $\approx$ are reserved for relations. The relation $\prec$ denotes the face relation between simplices of consecutive dimension.
2. Discrete Morse theory as generalized collapses
In this section, we revisit discrete Morse theory, framing it as a combinatorial generalization of collapsibility of simplicial complexes, in the context of Whitehead’s simple homotopy theory. The classical foundations of simple homotopy theory can be found in Whitehead’s seminal works [32, 33, 35], Milnor’s influential article [24], and Cohen’s textbook [13]. Throughout this manuscript, we will restrict our attention to finite simplicial complexes.
Definition 2.1.
Let $K$ and $L$ be simplicial complexes. There is an elementary collapse from $K$ to $L$ if there exist simplices $\sigma,\tau\in K$ such that $L=K\smallsetminus\{\sigma,\tau\}$ and $\tau$ is the unique simplex containing $\sigma$ properly in $K$. In this case, $\sigma$ is called a free face of $\tau$.
A collapse, denoted $K{\hskip 2.0pt\searrow\hskip 3.0pt}L$, is a sequence $K=K_{n}{\hskip 2.4pt\searrow^{e}\hskip 5.0pt}K_{n-1}{\hskip 2.4pt\searrow^{e}\hskip 5.0pt}\dots\hskip 2.4pt\searrow^{e}\hskip 5.0ptK_{0}=L$ of elementary collapses from $K$ to $L$. The inverse operation is called an expansion, denoted $L{\nearrow\hskip 3.0pt}K$. If $K{\hskip 2.0pt\searrow\hskip 3.0pt}L$ we say that $L$ is a weak core of $K$, and if $L$ has no free faces, we say it is minimal.
There is an $N$-deformation, denoted $K\hskip 2.0pt\diagup\searrow^{{}^{N}}\hskip 5.0ptL$, between $K$ and $L$ if there is a sequence $K=K_{n},K_{n-1},\dots,K_{0}=L$ such that $K_{i}{\hskip 2.0pt\searrow\hskip 3.0pt}K_{i-1}$ or $K_{i}{\nearrow\hskip 3.0pt}K_{i-1}$, and $\dim(K_{i})\leq N$ for all $i$. If there exists an $N$-deformation for some $N\in\mathbb{N}$ between $K$ and $L$, it is said that they have the same simple homotopy type.
Computing a minimal weak core of a simplicial complex provides a computational method for reducing its number of simplices while preserving its homotopy type. However, this approach is quite restrictive. For instance, many simplicial complexes lack free faces (e.g., triangulations of closed manifolds or famous examples such as the Dunce Hat [36] and Bing’s House [8]). Furthermore, a simplicial complex may have different (non-isomorphic) weak cores. For instance, if $D$ is any triangulation of the Dunce Hat, then $D\times I$ collapses to both $D$ and the singleton *. Thus, the order of collapses matters. A more general approach is provided by $N$-deformations (e.g., the above shows that $D\hskip 2.0pt\diagup\searrow^{3}\hskip 5.0pt*$, even though its minimal weak core is $D$ itself). Nonetheless, the existence of an $N$-deformation between two simplicial complexes not computable, as this would imply being able to algorithmically determine the contractibility of spaces.
Discrete Morse theory [18, 17], inspired by its smooth counterpart [23], provides a more general combinatorial framework for simplifying the structure of a simplicial complex while preserving its homotopy type. Here, deformations are encoded in terms of discrete Morse functions.
Definition 2.2.
Let $K$ be a simplicial complex. A function $f\colon K\to\mathbb{R}$ defined on the simplices of $K$ is called a discrete Morse function if, for every simplex $\sigma\in K$, the set
$M(\sigma)=\{\tau\prec\sigma\colon f(\tau)\geq f(\sigma)\}\cup\{\tau\succ\sigma\colon f(\tau)\leq f(\sigma)\}$
contains at most one element. A simplex $\sigma\in K$ is said to be critical if $M(\sigma)=\varnothing$.
To every simplicial complex $K$, we associate its face poset ${\mathcal{X}}(K)$, which is the poset of simplices of $K$ ordered by the face relation. Each Morse function determines a pairing of simplices $M=\big{\{}\{\sigma,\tau\}\colon M(\sigma)=\{\tau\}\big{\}}$. It can be shown that $M$ is a matching on the set of simplices of $K$. Moreover, the quotient of ${\mathcal{X}}(K)$ obtained by identifying matched simplices inherits a poset structure (see Section 4 for more details). This matching is referred to as an acyclic matching Chari [12] proved that acyclic matchings are in correspondence with discrete Morse functions.
The main theorem in discrete Morse theory establishes the connection between critical simplices of discrete Morse functions and the homotopy type of the simplicial complex.
Theorem 2.3 (Forman [15, 17, 18]).
Let $K$ be a finite simplicial complex of dimension $N$, and let $f\colon K\to{\mathbb{R}}$ be a discrete Morse function. Then, there is an $(N+1)$-deformation from $K$ to a CW-complex with exactly one cell of dimension $d$ for every critical $d$-simplex of $f$.
We denote by ${\mathrm{core}}_{f}(K)$ the CW-complex obtained from the critical simplices of $f$ in Theorem 2.3, and refer to it as the internal core of $K$ associated to $f$. It is important to note that Theorem 2.3 does not provide an explicit description of how the critical cells are attached along their boundaries in ${\mathrm{core}}_{f}(K)$. In [15, Prop. 1.3, Thm. 1.6], the author offers a recursive construction of the attaching maps for the critical cells, based on the finiteness of $K$; however, the construction is only made explicit for 2-dimensional cells.
The deformation from $K$ to ${\mathrm{core}}_{f}(K)$ is described locally by the Morse Lemma 2.4. Any discrete Morse function $f\colon K\to\mathbb{R}$ induces a filtration of $K$ by sublevel subcomplexes: for each $\alpha\in\mathbb{R}$, $K_{\alpha}$ is the smallest subcomplex of $K$ containing all simplices $\sigma$ with $f(\sigma)\leq\alpha$. The following result describes the evolution of the homotopy type complex along the filtration: intervals that do not contain critical simplices correspond to collapses, while those containing a single critical simplex account for homotopy changes.
Lemma 2.4 (Thm. 3.3 and 3.4, [17]).
Let $f\colon K\to\mathbb{R}$ be a discrete Morse function.
  1. A.
    If $f^{-1}\big{(}(\alpha,\beta]\big{)}$ contains no critical simplices, then $K_{\beta}{\hskip 2.0pt\searrow\hskip 3.0pt}K_{\alpha}$.
  2. B.
    If $f^{-1}\big{(}(\alpha,\beta]\big{)}$ contains exactly one critical $k$-simplex, then $K_{\beta}\simeq K_{\alpha}\cup e^{k}$, where $e^{k}$ is a $k$-dimensional cell.
From a combinatorial perspective, discrete Morse functions can be viewed as a totally ordered list of internal collapses.
Proposition 2.5 (Internal collapse [15]).
Let $K$ be a finite simplicial complex of dimension $N$. If a subcomplex $K_{0}\subseteq K$ collapses to a subcomplex $L_{0}$, then $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0ptL$, where $L$ is a CW-complex obtained from $L_{0}$ by attaching one $k$-cell for each $k$-simplex in $K\smallsetminus K_{0}$.
Proposition 2.5 also admits a formulation in terms of CW-complexes; see [15, Proposition 1.3]. We refer to the deformation from $K$ to $L$ as an internal collapse. The collapse takes place within a subcomplex $K_{0}\subseteq K$, while the resulting deformation affects the entire complex through the reattachment of cells corresponding to the simplices in $K\smallsetminus K_{0}$.
Internal collapses arise naturally in discrete Morse theory: for any discrete Morse function $f\colon K\to\mathbb{R}$, the transitions between sublevel complexes as described in the Morse Lemma (Lemma 2.4) correspond to internal collapses. Specifically, if the interval $(\alpha,\beta]$ contains no critical simplices, then $K_{\beta}\simeq K_{\alpha}$, and there is an internal collapse from $K$ to a CW-complex obtained from $K_{\alpha}$ by attaching a $k$-cell for each $k$-simplex in $K\smallsetminus K_{\beta}$. In general, the entire deformation from $K$ to ${\mathrm{core}}_{f}(K)$ can be interpreted as a well-ordered sequence of internal collapses.
We summarize the viewpoint of discrete Morse theory as a generalization of classical collapses in the following statement.
Proposition 2.6.
Let $K$ be a finite simplicial complex.
  1. (1)
    If $K\searrow L$, then there exists a discrete Morse function $f\colon K\to\mathbb{R}$ such that ${\mathrm{core}}_{f}(K)=L$.
  2. (2)
    If $f\colon K\to\mathbb{R}$ is a discrete Morse function, then there exists a sequence of internal collapses from $K$ to ${\mathrm{core}}_{f}(K)$.
3. Strong discrete Morse theory
Building on the strong homotopy theory for simplicial complexes introduced by Barmak and Minian [3], we present a version of discrete Morse theory centred on the concept of internal strong collapses. We recall first the notion of strong collapses.
Given a simplicial complex $K$ and a vertex $v\in V(K)$, the open star of $v$, denoted ${\mathrm{st}}(v,K)$, is the collection of simplices in $K$ that contain $v$. The subcomplex of simplices disjoint from $v$ is denoted $K\smallsetminus v$. The link of $v$ in $K$, denoted ${\mathrm{lk}}(v,K)$, is the subcomplex of $K\smallsetminus v$ consisting of simplices $\sigma$ such that $\sigma\cup\{v\}$ is a simplex of $K$. A simplicial cone over $K$ with apex $a$ (a vertex not in $K$) is the complex $a*K$ consisting of all simplices of $K$, the vertex $\{a\}$, and all simplices of the form $\sigma\cup\{a\}$ with $\sigma\in K$.
Definition 3.1 (Def. 2.1, [3]).
Let $K$ be a simplicial complex. A vertex $v$ of $K$ is dominated by a vertex $a$ if ${\mathrm{lk}}(v,K)$ is a simplicial cone $a*K_{0}$, where $K_{0}$ is a subcomplex of $K$. In that case, there is an elementary strong collapse from $K$ to $K\smallsetminus v$, denoted $K\searrow K\smallsetminus v$. A strong collapse $K\searrow\!\searrow L$ from a simplicial complex $K$ to a subcomplex $L$ is a sequence of elementary strong collapses starting at $K$ and ending at $L$. In that case, we say that $L$ is a strong core of $K$. If $L$ has no dominated vertices, we say that $L$ is a minimal strong core of $K$ [^2] This is simply called a core in [3]. We will use the term minimal strong core in this manuscript to avoid confusion with other core constructions.
If $v\in K$ is dominated by $a$, then there is an explicit strong deformation retraction $r_{a}\colon K\to K\smallsetminus v$ given by
$r_{a}(\sigma)=\begin{cases}\sigma&\text{if }v\notin\sigma,\\ \{a\}\cup\sigma\smallsetminus\{v\}&\text{if }v\in\sigma.\end{cases}$
More generally, if $K{\searrow\searrow\hskip 3.0pt}L$ via a sequence of elementary strong collapses
[…]
with $K_{i+1}=K_{i}\smallsetminus v_{i}$ and $v_{i}$ dominated by $a_{i}\in K_{i}$, then there is a strong deformation retraction $r\colon K\to L$ given by the composition
$r:=r_{a_{n}}\circ r_{a_{n-1}}\circ\dots\circ r_{a_{1}}.$
Moreover, if $K{\searrow\searrow\hskip 3.0pt}L$, then $K{\hskip 2.0pt\searrow\hskip 3.0pt}L$ [3, Rmk. 2.4]. That is, Whitehead’s notion of collapse is weaker than the notion of strong collapse.
Theorem 3.2 (Thm. 2.11. [3]).
Any sequence of strong collapses from a simplicial complex $K$ to a subcomplex without dominated vertices yields a unique (up to isomorphism) subcomplex of $K$. That is, the minimal strong core of a simplicial complex is unique up to isomorphism.
Recall that a CW-complex is regular if the attaching maps of all its cells are homeomorphisms onto their images. The following lemma formalizes the notion of an internal strong collapse, illustrated in Figure 2.
Lemma 3.3.
Let $K$ be a simplicial complex, and let $w$ be a vertex of $K$. Suppose there exists a subcomplex $L\subseteq K\smallsetminus w$ such that $K\smallsetminus w{\searrow\searrow\hskip 3.0pt}L$. Then $K$ is homotopy equivalent to a regular CW-complex
$Z:=L\cup\bigcup_{i}e_{i},$
where $L$ is embedded as a subcomplex of $Z$, and for each $k$-simplex $\sigma_{i}\in{\mathrm{st}}(w,K)$, there exists a corresponding $k$-cell $e_{i}$ not in $L$. The attaching map of each cell $e_{i}$ is induced from the attaching map of $\sigma_{i}$ and the strong deformation retraction $r\colon K\smallsetminus w\to L$ induced by the strong collapse.
If $\dim(K)=N$, then $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0ptZ$.
Proof.
Consider the pushout diagram
[…]
where $i$ is the inclusion and $r\colon K\smallsetminus w\to L$ is a strong deformation retraction induced by a strong collapse. By the universal property of the pushout, there exist induced maps $\bar{r}\colon K\to Z$ and $\bar{i}\colon L\to Z$ such that $\bar{r}\circ i=\bar{i}\circ r$. Since $r$ is a homotopy equivalence and $i$ is a closed cofibration, the Gluing Theorem (see, e.g., [11, 7.5.7, Corollary 2]) implies that $\bar{r}$ is also a homotopy equivalence.
The space $Z$ can be identified with the adjunction space $(K\sqcup L)/_{\sim}$, where $i(\sigma)\sim r(\sigma)$ for every simplex $\sigma\in K\smallsetminus w$. Let $q\colon K\sqcup L\to Z$ denote the corresponding quotient map. We define a CW-structure on $Z$ as follows. For each simplex $\sigma\in L$, we have $r(\sigma)=\sigma$, so $q$ acts as the identity and embeds $L$ into $Z$ with the same CW-structure it inherits from $K$. Moreover, since $r$ is simplicial, each cell $q(\sigma)$ for $\sigma\in K\smallsetminus w$ inherits a simplicial (hence regular) CW-structure from $K$. Now let $\sigma\in\operatorname{st}(w,K)$, and assume first that the strong collapse $K\smallsetminus w\searrow L$ is elementary, i.e., $L=(K\smallsetminus w)\smallsetminus v$ for some vertex $v$ dominated by $a\in K\smallsetminus w$. If $v\notin\sigma$, then $q(\sigma)=\sigma$ in $Z$, and the simplicial CW-structure is preserved. Similarly, if $v\in\sigma$ but $a\notin\sigma$, then $q(\sigma)=\{a\}\cup(\sigma\smallsetminus\{v\})$, which corresponds to a relabelling and again preserves regularity. The remaining case is when both $v$ and $a$ belong to $\sigma$. Suppose $\sigma=\{w,a,v_{0},\dots,v_{k}\}$, with $v_{0}=v$. Let $\bar{\sigma}\subseteq|K|$ denote the geometric realization of $\sigma$, which consists of formal convex combinations
$s\cdot a+r\cdot w+\sum_{i=0}^{k}t_{i}\cdot v_{i},\qquad\text{where }s+r+\sum t_{i}=1,\;s,r,t_{i}\geq 0.$
The quotient map $q$ identifies points in $\bar{\sigma}$ via the relation
$s\cdot a+\sum_{i=0}^{k}t_{i}\cdot v_{i}\approx(s+t_{0})\cdot a+\sum_{i=1}^{k}t_{i}\cdot v_{i},$
which encodes the strong collapse $K\smallsetminus w{\searrow\searrow\hskip 3.0pt}(K\smallsetminus w)\smallsetminus v$. that replaces the vertex $v_{0}=v$ by its dominator $a$ on the subcomplex The resulting quotient space $q(\bar{\sigma})$ corresponds to the internal collapse $w\ast{\mathrm{lk}}^{\downarrow}(v,K){\hskip 2.0pt\searrow\hskip 3.0pt}{\mathrm{lk}}^{\downarrow}(v,K)\subseteq\partial\bar{\sigma}$. Since we are identifying a contractible subcomplex with its image under a strong deformation retract, and this collapse occurs entirely within the boundary of $\bar{\sigma}$, the quotient $q(\bar{\sigma})$ remains homeomorphic to a closed $(k+2)$-ball $D^{k+2}$. Furthermore, the interior of $\bar{\sigma}$ is disjoint from the identified region and is mapped injectively to its image in $Z$. See [13, Prop. 3.4], [30, Ch. 4].
We define the corresponding cell $e_{\sigma}\subseteq Z$ to be the image of the interior of $\bar{\sigma}$, with closure $q(\bar{\sigma})\cong D^{k+2}$. The attaching map of $e_{\sigma}$ is then induced from the original attaching map $\varphi_{\sigma}\colon\partial\Delta^{k+2}\to K$ by composition with the quotient map. This yields a continuous map
$\bar{\varphi}_{\sigma}\colon\Delta^{k+2}/_{\approx}\;\to Z,$
where the quotient space $\Delta^{k+2}/_{\approx}$ is homeomorphic to $S^{k+1}$ and $\bar{\varphi}_{\sigma}$ is a homeomorphism onto its image. This defines a regular attaching map for the cell $e_{\sigma}$.
[…]
The general case follows by induction on the number of elementary strong collapses in the sequence
$K\smallsetminus w=K_{0}{\searrow\searrow^{e}\hskip 5.0pt}K_{1}\searrow\dots\searrow\searrow^{e}\hskip 5.0ptK_{n}=L.$
Inductively, we have a corresponding sequence of pushouts:
[…]
Here, $\bar{i}\colon K_{n-1}\to Z_{n-1}$ is an inclusion, and the same reasoning applied to $K$ extends to $Z_{n-1}$. By induction, the space $Z_{n-1}$ carries a regular CW-structure and is homotopy equivalent to $K$. $Z_{n}$ is obtained after performing an elementary strong collapse to $K_{n-1}$. To complete the inductive step, we must ensure that the new attaching maps are still a homeomorphism onto its image. The only scenario that could in principle compromise regularity would be if, for some simplex $\sigma\in{\mathrm{st}}(w,K)$, the entire boundary $\partial\sigma$ were identified to a single point in $Z_{n}$ due to prior identifications. However, this situation cannot arise: since $w$ is always a vertex of $\sigma$, and since strong collapses remove only dominated vertices, the vertex $w$ is never removed. Moreover, even if all other vertices of $\sigma$ are eventually collapsed, the image of the boundary of $\sigma$ must contain at least two distinct points — namely, $w$ and one vertex from $L$ not dominated during the collapse sequence. Hence, the attaching map remains a homeomorphism onto its image.
Therefore, the final space $Z$ is a regular CW-complex homotopy equivalent to $K$. The fact that this homotopy equivalence is an $(N+1)$-deformation follows from the observation that this process was obtained in particular by internal collapses, and by [15, Prop. 1.3], such collapses induce $(N+1)$-deformations. ∎
Proposition 3.4 (Strong internal collapse).
Let $K$ be a simplicial complex of dimension $N$. If a full subcomplex $K_{0}\subseteq K$ strong collapses to a subcomplex $L_{0}$, then $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0ptL$, where $L$ is a regular CW-complex obtained from $L_{0}$ by attaching one $k$-cell for each $k$-simplex in $K\smallsetminus K_{0}$.
Proof.
Follows by induction on the vertices in $V(K)\smallsetminus V(K_{0})$, using Lemma 3.3. ∎
Under the assumptions of Proposition 3.4, we say that there is a strong internal collapse from $K$ to $L$.
Let $K$ be a finite simplicial complex, and let $g\colon V(K)\to\mathbb{R}$ be a real-valued function on its vertex set. The function $f$ induces a filtration of $K$ by sublevel subcomplexes $\{K_{\alpha}\}_{\alpha\in{\mathbb{R}}}$, where each $K_{\alpha}$ is the subcomplex spanned by the vertices in the sublevel set $g^{-1}\big{(}(-\infty,\alpha]\big{)}$.
Definition 3.5.
Let $g\colon V(K)\to\mathbb{R}$ be a real-valued function, and let $v\in V(K)$ be a vertex. The descending open star of $v$, denoted ${\mathrm{st}}^{\downarrow}(v,K)$, is the open star of $v$ in the subcomplex $K_{g(v)}$. Similarly, the descending link of $v$, denoted ${\mathrm{lk}}^{\downarrow}(v,K)$, is the link of $v$ in $K_{g(v)}$. We say that $v$ is descending dominated if it is dominated in $K_{g(v)}$ by a vertex $a$ with $g(a)<g(v)$. Otherwise, $v$ is called a strong critical vertex.
We have an analogous version of the classical discrete Morse Lemma 2.4 in this setting.
Lemma 3.6.
Let $K$ be a simplicial complex and $g\colon V(K)\to\mathbb{R}$ a real-valued function.
  1. A.
    If $g^{-1}\big{(}(\alpha,\beta]\big{)}$ contains no strong critical vertices, then $K_{\beta}{\searrow\searrow\hskip 3.0pt}K_{\alpha}$.
  2. B.
    If $g^{-1}\big{(}(\alpha,\beta]\big{)}$ contains a single strong critical vertex $w$, then $K_{\beta}$ is homotopy equivalent to a CW-complex of the form
    $K_{\alpha}\cup\bigcup_{i}e_{i},$
    where the cells $e_{i}$ are in one-to-one correspondence with the simplices in ${\mathrm{st}}^{\downarrow}(w,K)$. This correspondence preserves dimensions, and the attaching maps of the cells are determined as in Lemma 3.3, resulting in a regular CW-complex structure.
Proof.
The function $g$ induces a total order $\{v_{1},v_{2},\dots,v_{n}\}$ on $V(K)$ such that $g(v_{i})\leq g(v_{j})$ whenever $i<j$. For each $i$, let $K_{i}$ be the subcomplex of $K$ generated by all simplices whose vertices lie in $\{v_{1},v_{2},\dots,v_{i}\}$.
For part A., assume that $g^{-1}\big{(}(\alpha,\beta]\big{)}$ contains no strong critical vertices. Then for each $v_{i}\in g^{-1}\big{(}(\alpha,\beta]\big{)}$, the descending link ${\mathrm{lk}}^{\downarrow}(v_{i},K)={\mathrm{lk}}(v_{i},K_{g(v_{i})})$ is a simplicial cone with apex $a_{i}$, where $g(a_{i})<g(v_{i})$. Since $K_{i}\subseteq K_{g(v_{i})}$, we have
${\mathrm{lk}}(v_{i},K_{i})={\mathrm{lk}}(v_{i},K_{g(v_{i})})\smallsetminus\{\sigma\in{\mathrm{lk}}(v_{i},K_{g(v_{i})})\colon\sigma\ni v_{j}\text{ with }j>i\text{ and }g(v_{j})=g(v_{i})\}.$
This subcomplex remains a cone with apex $a_{i}$. Therefore, $v_{i}$ is dominated in $K_{i}$ by $a_{i}$, and $K_{i}{\searrow\searrow^{e}\hskip 5.0pt}K_{i}\smallsetminus v_{i}=K_{i-1}$. Applying this inductively, we obtain a sequence of strong collapses from $K_{\beta}$ to $K_{\alpha}$.
Part B. follows from part A. and Lemma 3.3. ∎
Building on Proposition 3.4 and Lemma 3.6, we establish the following result, which formalize the relationship between internal strong collapses, regular CW-complexes, and discrete Morse theory.
Theorem 3.7 (Strong Discrete Morse Theory).
Let $K$ be a finite simplicial complex, and let $g\colon V(K)\to\mathbb{R}$ be a real-valued function. Then, $K$ is homotopy equivalent to a regular CW-complex ${\mathrm{core}}_{g}(K)$, whose cells are in one-to-one correspondence with the simplices in the descending open star of the strong critical vertices of $g$.
Proof.
Subdivide the image of $g$ into subintervals, each satisfying the conditions of Lemma 3.6 A. or B., and iteratively apply the corresponding deformations to $K$. The cells in the resulting CW-complex correspond to the simplices in $\{{\mathrm{st}}^{\downarrow}(w,K)\colon w\text{ is a critical vertex in }K\}$. ∎
We denote by ${\mathrm{core}}_{g}(K)$ the regular CW-complex obtained from the critical simplices of $g\colon V(K)\to\mathbb{R}$ in Theorem 3.7, and refer to it as the strong internal core of $K$ associated to $g$. Notice that the strong internal core depends on a choice of dominating vertices for the strong internal collapses, but we omit this information unless necessary.
Proposition 3.8 (Connection to Strong Collapses).
If $K{\searrow\searrow\hskip 3.0pt}L$, then there exists a function $g\colon V(K)\to\mathbb{R}$ such that ${\mathrm{core}}_{g}(K)=L$.
Proof.
Let $v_{1},v_{2},\dots,v_{n}\in V(K)$ be the sequence of dominated vertices corresponding to a strong collapse from $K$ to $L$, i.e., define $K_{n}=K$ and $K_{i-1}=K_{i}\smallsetminus v_{i}$, with each $K_{i}{\searrow\searrow^{e}\hskip 5.0pt}K_{i-1}$ and $K_{0}=L$. Define a function $g\colon V(K)\to\mathbb{R}$ by
$g(v)=\begin{cases}i&\text{if }v=v_{i},\\ 0&\text{if }v\in V(L).\end{cases}$
We claim that the strong internal core, ${\mathrm{core}}_{g}(K)$, coincides with $L$. By construction, each $v_{i}$ is dominated in $K_{i}$ by some vertex $a_{i}\in K_{i}$. Since $v_{i}$ is removed at stage $i$, and $a_{i}\in K_{j}$ for some $j<i$, it follows that $g(a_{i})<g(v_{i})$. Therefore, $v_{i}$ is descending dominated. On the other hand, any vertex $v\in V(L)$ satisfies $g(v)=0$, and hence $g^{-1}((-\infty,g(v)))=\varnothing$. Thus, $v$ cannot be descending dominated and is a strong critical vertex.
Hence, the only strong critical vertices of $g$ are those in $L$, and the strong internal core, ${\mathrm{core}}_{g}(K)$, coincides with $L$. ∎
Theorem 3.9 (Connection to discrete Morse theory).
Let $K$ be a finite simplicial complex and let $g\colon V(K)\to{\mathbb{R}}$ be a real-valued map. Then, there exists a discrete Morse function $f\colon K\to{\mathbb{R}}$ on $K$ such that its critical simplices are the simplices in {st(w,K):w is a strong critical vertex
 of
g
[…]
 of
[…]
. Moreover, ${\mathrm{core}}_{g}(K)\cong{\mathrm{core}}_{f}(K)$, that is the strong internal core associated to $g$ is precisely the internal core associated to $f$.
Proof.
Let $C=\{{\mathrm{st}}^{\downarrow}(w,K):w\text{ is a strong critical vertex of }g\}\subseteq K$. Given the correspondence between discrete Morse functions and acyclic matchings [12], we will construct an acyclic matching $M$ with critical simplices $C$ as follows. For each simplex $\sigma\in K$ containing descending dominated vertices, let
$v_{\sigma}=\arg\max\{f(v)\colon v\in\sigma\text{ is a descending dominated vertex}\},$
and let $a_{\sigma}$ denote the vertex dominating $v_{\sigma}$. The pairing $M$ is defined by
$(\sigma\smallsetminus\{a_{\sigma}\},\sigma)\in M\text{ if }a_{\sigma}\in\sigma,\quad\text{or}\quad(\sigma,\{a_{\sigma}\}\cup\sigma)\in M\text{ if }a_{\sigma}\notin\sigma.$
To verify that $M$ is a matching, assume for contradiction that a simplex in $K$ is matched to more than one simplex. Without loss of generality, suppose $(\sigma,\{a\}\cup\sigma)\in M$ with $v_{\sigma}$ the associated descending dominated vertex. If there exists another pair ($(\{a\}\cup\sigma,\{a^{\prime},a\}\cup\sigma)\in M$, then $a^{\prime}$ dominates a vertex $v^{\prime}\in\sigma$. Since $v^{\prime}$ must also satisfy $v^{\prime}=\arg\max\{f(v)\colon v\in\sigma$ is descending dominated}}\}} this contradicts the uniqueness of $v_{\sigma}$. Alternatively, if there exists a pair $(\sigma^{\prime},\{a\}\cup\sigma)\in M$, then $a\sigma=\{a^{\prime}\}\cup\sigma^{\prime}$ implies $\sigma=\{a^{\prime}\}\cup\tau$ and $\sigma^{\prime}=\{a\}\cup\tau$ for some simplex $\tau$. This forces $\tau$ to contain two distinct dominated vertices, $v_{\sigma}$ and $v^{\prime}$, each maximizing $f(v)$, which is again a contradiction. Therefore, $M$ is a valid matching.
To establish that $M$ is acyclic, consider any pair $(\sigma,\{a\}\cup\sigma)\in M$. For any $\tau\succ\sigma$, we claim that $(\tau\smallsetminus\{a^{\prime}\},\tau)\notin M$ for any $a^{\prime}\neq a$. Since $\tau\succ\sigma$, the simplex $\tau$ contains $v_{\sigma}$. If $v_{\sigma}$ also satisfies $v_{\sigma}=\arg\max\{f(v)\colon v\in\tau\text{ is a descending dominated vertex}\}$, then $(\tau,\{a\}\cup\tau)\in M$, as $a$ dominates $v_{\sigma}$. Alternatively, if there exists another dominated vertex $w\in\tau$ such that $w>v_{\sigma}$, then $(\tau,\{a_{w}\}\cup\tau)\in M$, where $a_{w}$ dominates $w$. In either case, $(\tau\smallsetminus\{a^{\prime}\},\tau)\notin M$ for any $a^{\prime}\neq a$, ensuring that $M$ is acyclic. This completes the proof. ∎
4. The construction of the internal core
Given a CW-complex $K$, its face poset ${\mathcal{X}}(K)$ encodes the incidence relations among its cells: $e\leq e^{\prime}$ in ${\mathcal{X}}(K)$ if $\overline{e}\subseteq\overline{e^{\prime}}$. While this poset does not determine the homotopy type of $K$ in general, in the case of regular CW-complexes it serves as a complete combinatorial model: the order complex [^3] The order complex of a finite poset $X$ is the simplicial complex whose simplices are the non-empty totally ordered subsets of $X$. of ${\mathcal{X}}(K)$, denoted ${\mathcal{K}}({\mathcal{X}}(K))$, is homotopy equivalent to $K$ (see [9]).
Let $X$ be a finite poset. An acyclic matching on $X$ is a collection $M$ of disjoint pairs $\{x,y\}\subseteq X$ such that the quotient set $X/_{\sim}$, where $x\sim y$ if $\{x,y\}\in M$, admits a partial order induced by the original order on $X$ (the relation $[x]\leq[y]\in X/_{\sim}$ if $x\leq y$ in $X$ is antisymmetric).
Discrete Morse functions on a simplicial complex $K$ are in bijection with acyclic matchings on ${\mathcal{X}}(K)$, where the unmatched elements correspond to the critical simplices [12]. Given an acyclic matching $M$ on ${\mathcal{X}}(K)$, one can construct auxiliary posets that reflect the structure of the Morse reduced CW-complex. Inspired by Nanda’s work on discrete Morse theory and localization of categories [25], we introduce the following definitions.
Definition 4.1.
Let $K$ be a simplicial complex, and let $M$ be an acyclic matching on ${\mathcal{X}}(K)$. The localization poset ${\mathrm{Loc}}_{M}(K)$ is the poset obtained by identifying matched simplices in ${\mathcal{X}}(K)$. The critical poset ${\mathrm{Crit}}_{M}(K)$ is the subposet of ${\mathrm{Loc}}_{M}(K)$ consisting of equivalence classes of unmatched (i.e., critical) simplices.
In general, neither ${\mathrm{Crit}}_{M}(K)$ nor ${\mathrm{Loc}}_{M}(K)$ determine the homotopy type of $K$, even when $K$ is regular or simplicial (see Example 4.3). However, when $M$ arises from internal strong collapses, Theorem 3.9 implies that ${\mathrm{Crit}}_{M}(K)$ is a finite model of $K$.
Corollary 4.2.
Let $K$ be a simplicial complex and let $g\colon V(K)\to\mathbb{R}$ be a real-valued function. Then, there exists an acyclic matching $M_{g}$ on ${\mathcal{X}}(K)$ such that the critical poset ${\mathrm{Crit}}_{M_{g}}(K)$ is the face poset of ${\mathrm{core}}_{g}(K)$. In particular, ${\mathrm{Crit}}_{M_{g}}(K)$ is a finite model of $K$.
The critical poset can be constructed algorithmically, as detailed in Algorithm 1.
Algorithm 1 The critical poset
1:Face poset $\mathcal{X}(K)$ of a simplicial complex $K$, set of critical simplices $C$, acyclic matching $M$ on $\mathcal{X}(K)$
2:Critical poset ${\mathrm{Crit}}_{M}(K)$
3:$P\leftarrow$ elements of $\mathcal{X}(K)$
4:$R\leftarrow$ cover relations of $\mathcal{X}(K)$
5:Define a map ${\mathrm{QuotientClass}}:P\to\text{new elements}$
6:for all $(\sigma,\tau)\in M$ do
7:     $P\leftarrow P\setminus\{\sigma,\tau\}$
8:     $[\sigma\sim\tau]\leftarrow$ new element representing $\{\sigma,\tau\}$
9:     $P\leftarrow P\cup\{[\sigma\sim\tau]\}$
10:     ${\mathrm{QuotientClass}}[\sigma]\leftarrow[\sigma\sim\tau]$
11:     ${\mathrm{QuotientClass}}[\tau]\leftarrow[\sigma\sim\tau]$
12:end for
13:for all $\sigma\in P$ such that ${\mathrm{QuotientClass}}[\sigma]$ is undefined do
14:     ${\mathrm{QuotientClass}}[\sigma]\leftarrow\sigma$
15:end for
16:$R^{\prime}\leftarrow\emptyset$
17:for all $(\sigma,\tau)\in R$ do
18:     if ${\mathrm{QuotientClass}}[\sigma]\neq{\mathrm{QuotientClass}}[\tau]$ then
19:         $R^{\prime}\leftarrow R^{\prime}\cup\big{(}{\mathrm{QuotientClass}}[\sigma],{\mathrm{QuotientClass}}[\tau]\big{)}$
20:     end if
21:end for
22:${\mathrm{Loc}}_{M}(K)\leftarrow$ poset with elements $P$ and relations $R^{\prime}$
23:${\mathrm{Crit}}_{M}(K)\leftarrow$ subposet of ${\mathrm{Loc}}_{M}(K)$ induced by $C$
24:return ${\mathrm{Crit}}_{M}(K)$
Example 4.3.
Let $K$ be the boundary of the 3-simplex. Consider the acyclic matching depicted in Figures 3 and 4. The posets ${\mathrm{Loc}}_{M}(K)$ and ${\mathrm{Crit}}_{M}(K)$ are contractible, and do not reflect the homotopy type of $K\simeq S^{2}$.
Figure 3. The boundary of a 3-simplex endowed with a function $g\colon V(K)\to\mathbb{R}$ on vertices, together with an acyclic matching $M$ on the simplices of $K$. Red arrows indicate matched pairs of simplices, oriented from lower- to higher-dimensional ones. By abuse of notation, and due to the injectivity of $g$, we identify each vertex with its image under $g$.
${\bullet_{(0,1,2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,1,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,2,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(1,2,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,1)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(1,2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(1,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(2,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0)}}$${\bullet_{(1)}}$${\bullet_{(2)}}$${\bullet_{(3)}}$${{\mathcal{X}}(K)}$
(0,1,2)[(0,2),(0,2,3)][(0,3),(0,1,3)][(0),(0,1)][(1,3),(1,2,3)][(1),(1,2)][(2),(2,3)](3)${{\mathrm{Loc}}_{M}(K)}$
   (0,1,2)(3)${{\mathrm{Crit}}_{M}(K)}$
Figure 4. The face poset ${\mathcal{X}}(K)$, the localization poset ${\mathrm{Loc}}_{M}(K)$ and the critical poset ${\mathrm{Crit}}_{M}(K)$ associated to the acyclic matching $M$ in red (c.f. Fig 3).
On the other hand, consider the function $g\colon V(K)\to{\mathbb{R}}$ on the vertices depicted in Figure 3. For the associated acyclic matching $M$, the critical poset ${\mathrm{Crit}}_{M}(K)$ recovers the homotopy type of $K$, and it is the face poset of the strong internal core (see Figure 5).
${\mathrm{core}}_{g}(K)$
${\bullet_{(0,1,2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,1,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,2,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(1,2,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,1)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(1,2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(1,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(2,3)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bullet_{(0)}}$${\bullet_{(1)}}$${\bullet_{(2)}}$${\bullet_{(3)}}$${{\mathcal{X}}(K)}$
[…]
Figure 5. The strong internal core associated to the map $g\colon V(K)\to{\mathbb{R}}$ on vertices (Fig. 3), the face poset ${\mathcal{X}}(K)$ (with the acyclic matching $M_{g}$ induced by internal strong collapses, in red) and the critical poset ${\mathrm{Crit}}_{M_{g}}(K)$. Here, ${\mathrm{core}}_{g}(K)$ has two 0-cells, three 1-cells and three 2-cells.
Remark 4.4.
Consider the order-preserving maps
${\mathcal{Q}}\colon{\mathcal{X}}(K)\to{\mathrm{Loc}}_{M}(K)\quad\text{and}\quad{\mathcal{J}}\colon{\mathrm{Crit}}_{M}(K)\to{\mathrm{Loc}}_{M}(K),$
where ${\mathcal{Q}}$ denotes the quotient map and ${\mathcal{J}}$ the inclusion map. In future work, we will establish general conditions on the acyclic matching $M$ under which these maps induce homotopy equivalences at the level of simplicial complexes. In particular, we will show that when $M$ is induced by internal strong collapses, both ${\mathcal{Q}}$ and ${\mathcal{J}}$ induce homotopy equivalences.
5. Algorithmic simplification of simplicial complexes
Let $K$ be a simplicial complex of dimension $N$. We summarize below the different core notions associated to $K$:
  • Strong core: A subcomplex $L\subseteq K$ such that $K{\searrow\searrow\hskip 3.0pt}L$. The minimal strong core (i.e. with no dominated vertices) is unique up to isomorphism.
  • Weak core: A subcomplex $L\subseteq K$ such that $K{\hskip 2.0pt\searrow\hskip 3.0pt}L$. A minimal weak core (i.e. with no free faces) is not necessarily unique.
  • Internal core: Given a discrete Morse function $f\colon K\to\mathbb{R}$, the internal core ${\mathrm{core}}_{f}(K)$ is the CW-complex obtained by performing the sequence of internal collapses determined by $f$. It satisfies $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0pt{\mathrm{core}}_{f}(K)$. The internal core is not necessarily regular.
  • Strong internal core: Given a function $g\colon V(K)\to\mathbb{R}$ defined on the vertices of $K$, the strong internal core, ${\mathrm{core}}_{g}(K)$, is obtained by the sequence of internal strong collapses guided by $g$. It satisfies $K\hskip 2.0pt\diagup\searrow^{{}^{N+1}}\hskip 5.0pt{\mathrm{core}}_{g}(K)$, and the resulting space is a regular CW-complex.
The diagram below shows implication relations among the core constructions.
Internal CoreStrong Internal CoreWeak CoreStrong CoreThm. 3.9Prop. 2.6Prop. 3.8[3, Rmk 2.4]
For any method of reduction of simplicial complexes, computing the associated core requires the choice of an ordered sequence of collapses. While the minimal strong core is unique (up to isomorphism) regardless of the sequence of strong collapses, a minimal weak core does depend on the chosen sequence. Similarly, the internal core of a simplicial complex depends on the discrete Morse function. Note that (strong) discrete Morse theory can be seen as a systematic procedure for performing a sequence of (strong) reductions. These reductions may consist of (strong) internal collapses, or, when no free face (resp. dominated vertex) is available, the removal of a maximal simplex (resp. vertex). To address the dependence on the choice of the reduction sequence, one approach is to explore all possible sequences. However, this can be computationally prohibitive due to the exponential growth in the number of options. Alternatively, random reductions — where reductions are chosen randomly among the available options at each step — provide an efficient strategy to deal with this dependence.
To compute the minimal strong internal core, we iteratively apply a greedy algorithm that performs randomly chosen internal strong collapses whenever a dominated vertex is available. If no dominated vertex exists, we randomly select a vertex to be removed. Algorithm 2 formalizes this procedure. As part of this procedure, we construct the acyclic matching associated to the internal strong collapses as stated in Theorem 3.9. Notice that Algorithm 2 can also be used to study $(N+1)$-deformations, as the strong internal core $(N+1)$-deforms to the original complex. A similar procedure can be applied for the randomized computation of a minimal weak core of a simplicial complex.
Remark 5.1.
The worst-case complexity of a naive implementation of Algorithm 2 is $O(V^{3}\times F)$, where $V$ is the number of vertices of the simplicial complex $K$, and $F$ is the number of simplices of $K$.
Algorithm 2 A random strong internal core
${\mathcal{X}}(K)$: the face poset of a simplicial complex $K$
The critical poset of a random strong internal core of $K$
function StrongMorseReduction($X,C,M$)
$\triangleright$ Auxiliary function. Input: $X$: face poset, $M$: acyclic matching, $C$: critical elements
     if $X=\emptyset$ then
         return $(X,C,M)$
     end if
     $S\leftarrow\text{Shuffle}(\text{MinimalElements}(X))$triangleright Randomized linear extension of minimal elements
     for $v\in S$ do
         if there exists $a\in\text{MinimalElements}(X)\setminus\{v\}$ such that $a$ dominates $v$ then
              for all $\sigma\in\{\sigma\in X\colon v\leq\sigma\}$ such that $a\nleq\sigma$ do
                  $M\leftarrow M\cup\{(\sigma,\{a\}\cup\sigma)\}$
              end for
              $X\leftarrow X\setminus\{\tau\in X\colon v\leq\tau\}$
              return StrongMorseReduction($X,C,M$)
         end if
     end for
     $v\leftarrow\text{First}(S)$triangleright Take first non-dominated minimal element
     $C\leftarrow C\cup\{\sigma\in X\colon v\leq\sigma\}$
     $X\leftarrow X\setminus\{\sigma\in X\colon v\leq\sigma\}$
     return StrongMorseReduction($X,C,M$)
end function
procedure ComputeInternalStrongCore(${\mathcal{X}}(K)$)
     $(-,C,M)\leftarrow$ StrongMorseReduction(${\mathcal{X}}(K),\emptyset,\emptyset$)
     return CriticalPoset(${\mathcal{X}}(K),C,M$)
end procedure
In general, there is no combinatorial description of the internal core determined by a general discrete Morse function. Benedetti and Lutz [7] performed a computational analysis of the topology of simplicial complexes using randomized discrete Morse theory, studying the number of critical cells arising from random discrete Morse functions.
We analyse the examples from the Library of Triangulations [6] and compare the results of minimals weak core, strong core, and strong internal core, cases where the reduced complex can be completely reconstructed. We perform 10 iterations of the computation of random simple and strong internal cores for the simplicial complexes listed in Table 1, and 100 iterations for those in Table 2. Additionally, we implement a combined reduction strategy, where we first compute a random minimal weak core of the original complex and subsequently compute a random strong internal core of the resulting complex. Tables 1 and 2 present the mean sizes of the different cores and the corresponding computation times [^4] The experiments were conducted on a MacBook Pro with an Apple M3 Pro chip, 12-core CPU, 18-core GPU, and 18 GB of unified memory.. The implementation was developed in SageMath [14], and the code is available at https://github.com/ximenafernandez/Strong-Morse-Theory.
We observe that in cases where there are neither dominated vertices nor free faces, internal strong collapses allow significant reductions in the size of the complex, with reductions of 30% to 40% in instances as BH_3, BH_4, Bernadette_Sphere, and non_PL and 50% to 60% in the large examples trefoil_bsd and Hom_C5_K4. Notably, in many of these cases, the vector of critical simplices in a (classical) discrete Morse function does not fully determine the homotopy type. For example, the 18-vertex non-PL triangulation of the 5-dimensional sphere $S^{5}$ (non_PL) admits $(1,0,0,2,2,1)$ as its smallest discrete Morse vector (see [7]). In several large cases with available simple collapses, the strong Morse theory approach yields a more efficient model, significantly reducing the complex size by around 80% and computation times by over 90% in examples such as knot and bing.
On the other hand, a few cases exhibit simplicial complexes with small weak cores but large strong internal cores (on average). Examples include B_3_9_18, trefoil_arc, rudin, and triple_trefoil_arc. In these scenarios, the combined strategy emerges as the optimal choice: preprocessing the simplicial complex by computing its minimal weak core before determining the strong internal core.
Name example Number of simplices/Time per iteration (sec)
Original Minimal strong core Minimal weak core Strong internal core Strong internal core of minimal weak core
bing 8131 2259 2377.6 1665.80 1581.20
185.7559 1532.325 112.5924 1586.5839
non_4_2_colorable 5982 5982 5982.00 5838.20 5857.40
2.2379 1.5879 78.1108 79.2957
Hom_C5_K4 6240 6240 6240.00 4632.40 4669.60
3.0053 1.6739 139.9162 141.8681
trefoil_bsd 5876 5876 5876.00 2072.20 2134.80
3.7306 1.5139 102.2188 110.9461
knot 6203 1639 1946.20 1184.00 1271.60
86.4477 2088.4905 56.3632 2130.4097
Table 1. Comparison of reduction methods for simplicial complexes, averaged over 10 iterations. Bold values indicate cases where strong internal collapses yield the most significant size reduction. Blue values indicate cases where strong internal collapses also lead to a significant improvement in execution time compared to other methods.
Name example Number of simplices/Time per iteration (sec)
Original Minimal Strong core Minimal weak core Strong internal core Strong internal core of minimal weak core
Abalone 101 101 101.00 73.30 70.58
0.0273 0.0009 0.0484 0.0486
BH 131 131 131.00 100.68 99.60
0.0185 0.0015 0.0839 0.0824
BH_3 301 301 301.00 206.32 207.78
0.0401 0.0068 0.4438 0.4473
BH_4 401 401 401.00 277.06 274.60
0.0564 0.0124 0.8285 0.8371
BH_5 501 501 501.00 349.08 343.86
0.0806 0.0197 1.4011 1.4009
d2_n8_3torsion 49 49 49.00 38.48 38.38
0.0060 0.0003 0.0135 0.0138
d2_n8_4torsion 53 53 53.00 43.66 43.98
0.0060 0.0003 0.0135 0.0138
d2_n9_5torsion 65 65 65.00 57.12 55.68
0.0073 0.0005 0.0193 0.0204
dunce_hat 49 49 49.00 37.60 38.84
0.0056 0.0003 0.0135 0.0129
d2n12g6 122 122 122.00 118.80 118.66
0.0140 0.0013 0.00602 0.0612
regular_2_21_23_1 266 266 266.00 259.52 259.52
0.0333 0.0058 0.2230 0.2349
rand2_n25_p0.328 1076 1076 1074.00 1069.90 1067.18
0.1965 0.2955 1.3532 1.6397
dunce_hat_in_3_ball 75 1 1.00 1.00 1.00
0.0149 0.0151 0.0119 0.0153
Barnette_sphere 92 92 92.00 40.66 39.58
0.0130 0.0005 0.0182 0.0195
B_3_9_18 103 59 1.00 31.38 1.00
0.0220 0.0286 0.0212 0.0287
trefoil_arc 193 193 1.76 148.68 1.62
0.0281 0.1257 0.0865 0.1262
trefoil 250 250 250.00 206.98 211.12
0.0369 0.0031 0.1238 0.1298
rudin 215 215 1.00 137.74 1.00
0.0329 0.1668 0.1057 0.1670
poincare 392 392 392.00 361.80 364.24
0.0631 0.0074 0.3636 0.2744
double_trefoil 400 400 400.00 380.20 383.08
0.0643 0.0078 0.2714 0.2797
triple_trefoil_arc 449 449 172.36 438.62 157.60
0.0746 1.3937 0.3337 1.5215
triple_trefoil 536 536 536.00 526.50 525.70
0.0922 0.0141 0.4367 0.4467
hyperbolic_dodecahedral_space 718 718 718.00 708.82 709.36
0.1313 0.0241 23.3319 74.4473
S_3_50_1033 4232 4232 4232.00 4198.36 4295.20
1.5958 0.8286 16.9282 17.73918
600_cell 2640 2640 2640.00 2343.46 2349.68
0.6912 0.3057 16.5970 16.9279
CP2 255 255 255.00 219.78 219.68
0.0464 0.0020 0.0798 0.0823
RP4 991 991 991.00 942.36 937.34
0.2400 0.0293 0.7976 0.8290
K3_16 1704 1704 1704.00 1691.34 1690.62
0.4895 0.09545 1.61122 1.7076
K3_17 1854 1854 1854.00 1835.32 1836.22
0.6331 0.1108 1.9267 2.0365
RP4_K3_17 2813 2813 2813.00 2757.98 2751.66
0.9971 0.2429 5.2140 5.4425
RP4_11S2xS2 3179 3179 3179.00 3099.04 3105.62
1.2477 0.3071 6.8191 7.148
SU2_SO3 1534 1534 1534.00 1474.70 1471.84
0.5173 0.0453 1.2220 1.2677
non_PL 2608 2608 2608.0 1690.56 1539.16
1.1274 0.1274 2.8094 2.6714
Table 2. Comparison of algorithmic reduction methods for simplicial complexes, averaged over 100 iterations. Bold values highlight cases where significant reduction is achieved via strong internal collapses (compared to other methods). Red values indicate cases where significant reduction is only achieved through Whitehead collapses.
Acknowledgments
The author thanks Robert Green for his enthusiasm for posets and for pointing out the earlier work of Nicholas Scoville on Strong Morse Theory during his visit to Oxford. This work was inspired by fruitful conversations with him at the Mathematical Institute. The author is also grateful to Vidit Nanda for engaging with early versions of this project and for suggesting the connection with the localization of categories.